OpenStudy (anonymous):

y' + y*x^2 = 0 y(0) = -2

OpenStudy (anonymous):

\[\frac{dy}{dx} + y*x^2 = 0 \] \[ \frac{dy}{dx} = -y*x^2\] \[ \frac{1}{y}dy = -x^2dx\] \[\log_e y + C_1 =-\frac{x^3}{3} + C_2\] \[ \log_ey +\frac{x^3}{3} + C = 0 \] \[C = C_1 - C_2\]

OpenStudy (anonymous):

y(x) = c_1 e^(-x^3/3)